To have a status of the 50Hz in the PSTAB photodiodes I have made some plots using the time when the master laser was swapped. This provided a time without a beam at 12:00 UTC on May 11, and time with the PSTAB loop closed on May 12 at 00:00 UTC.
Figure 1 is an amplitude spectral density (V/rtHz) and Figure an amplitude spectrum (V) with a 10s duration FFT. Figure 3 and 4 are the same data using 1s duration FFT. For the ASD the broadband noise level doesn't depend on the resolution, while for the AS the 50Hz line height does not depend on the resolution, as should be the case.
Note that the dark noise spectra are higher than when the loop is closed. As the photodiode is in loop, and the high gain of the loop suppress the noise below the PD sensing noise. I couldn't find the out-of-loop photodiode PSTAB_PD1_AC channel in the data, it was present during O3.